# Testing the Origin of Hot Jupiters with Ariel

### v1

@@ 5.3 Suggestion for Ariel
Under a uniform period distribution simulating 250 planets, separated into 29 period bins, we find that Ariel observations at the Tier 2 level of uncertainties are sufficient to test the effect of high-eccentricity migration of hot Jupiters by way of their elevated water abundance ratio compared to warm Jupiters (Figure 7). Even in the relatively pessimistic scenario (perfect dust trap; left panel), the simulated 250 planets having 8 data points within each bin appear more than enough to beat down the stochastic noise and extract the 0.5 dex difference in H2O abundance ratio between hot and warm Jupiters at 1-. We therefore suggest having at minimum 100 targets to test the effect of high-eccentricity migration of hot Jupiters in comparison to warm Jupiters.
Potentially, a bigger challenge is to sample enough hot and warm Jupiters. Figure 6 shows that Tier 2 MCS sample we use is highly concentrated around 2–5 days of orbital period. Although this orbital period corresponds to the peak of the orbital period distribution of hot Jupiters, the distribution extends to 10 days (e.g., A. Santerne et al., 2016; S. W. Yee & J. N. Winn, 2023).555While Figure 7 shows a clear jump in water abundance at 4 days, the location of the jump is arbitrary as it is completely set by our choice of transporting planets down to 1–4 days to simulate high-eccentricity migration. It is therefore important to use an independent definition of hot vs. warm Jupiters which are usually divided into those at orbital periods less than and greater than 10 days based on occurrence rate studies (e.g., A. Santerne et al., 2016; S. W. Yee & J. N. Winn, 2023). Based on our filtered MCS, we find that the number of warm Jupiters (orbital periods 10 days) for which we can obtain Tier 2 spectra within the first 100 days is only 7. This number increases to 11 within the first 150 days (6 inside 15 days, 5 beyond) and to 16 within a full year (all additional 5 targets with orbital periods inside 15 days). This number is more than sufficient to reach the target precision to distinguish between hot and warm Jupiter water abundance under high-eccentricity migration of the former if we group all warm Jupiters under one period bin; it is marginally sufficient if we group warm Jupiters over two period bins, separated at 15 days.
@@ 6 Conclusion
We explored three post-formation pollution mechanisms—planetesimal accretion, pebble accretion, and disk-induced migration—to construct a theoretical prediction of gas giants’ atmospheric water abundance as a function of orbital period. Only pebble accretion in highly metal-heavy disks is able to achieve supersolar metallicity in giant planet atmosphere leaving a detectable signature of pollution. If the dominant origin channel of hot Jupiters is high-eccentricity migration while that of warm Jupiters is more dynamically quiescent, we expect a significant upward shift in their atmospheric water abundance compared to warm Jupiters, at a level that is detectable with Ariel. To ensure that it can constrain these formation/migration scenarios, the Ariel Tier 2 transit survey should include hot (10 days) and warm (10 days) Jupiters with at least 8–16 warm Jupiters beyond 10 days. We further suggest a comparison of median hot Jupiter water abundance with that of wide-orbit cold Jupiters which can be leveraged to reveal the property of primordial protoplanetary disks such as the existence and the efficacy of the dust trap established by the embedded giant planet(s).

### v3

@@ 5.1 Stellar and Planet Metallicity
Further astrophysical scatter can be introduced by any metallicity dependence on planet mass or disk solid mass reservoir (e.g., D. P. Thorngren et al., 2016). Under our pebble accretion scenario, the accreted solid mass is much more strongly dependent on the disk solid surface density than the planet mass (equation 16) so we expect the scatter introduced by the disk mass to be greater. Alternatively, a scatter in planet metallicity can be achieved by giant impact between gas giants of varying core and gas masses (S. Ginzburg & E. Chiang, 2020). Since the exact source of a mass-metallicity relation or its scatter remains unclear, we compute the required number of targets so that we can still discern the differences in water abundances between hot (10 days) and warm (10 days) Jupiters at 2–3 level for a varying additional scatter in the metallicity data (i.e., in this test, we only have two bins in orbital periods, separated at 10 days).
The result of our analysis for pebble accretion with dust trapping is shown in Figure 9. For 2- detections, the required number of targets is below the 232 expected total number of targets that can reach Tier 2 signal-to-noise within 1 year even if the additional astrophysical scatter is 1.0 dex. In the case of 3- detections, more targets are required with additional scatter 0.2 dex. However, with more targets the observation time exceeds the allotted mission time after just 0.4 dex of additional scatter, since the remaining warm Jupiters require far more time to reach Tier 2. This is a limitation of the dataset, as there are only 134 warm Jupiters with under 50 day periods, many of which are poor Tier 2 transmission targets. Although having more warm Jupiters would inevitably improve the results, it is expected that the final Ariel target sample will be weighted towards the hot Jupiters as they are more easily characterizable and more numerous, perhaps more aggressively than the ratio used here.
@@ 5.2 Formation Condition of Hot Jupiters
Classically, planetesimal pollution has been invoked to explain the high atmospheric metallicity of Jovian planets in solar (e.g., O. Mousis et al., 2009) and extrasolar systems (e.g., J. J. Fortney et al., 2013; C. Mordasini et al., 2016). Our calculations disfavor planetesimal accretion on the basis of the conservative limit of the accretion cross-sectional area presented by the planet’s geometric cross section. If the random velocities of the planetesimals can be sufficiently damped below the surface escape velocity of the planets, gravitational focusing will enhance the rate of planetesimal accretion (see P. Goldreich et al., 2004, for a review) and at its maximal capacity where the accretion radius scales with the Hill radius of the planet, the amount of accreted solid would be comparable to what we expect from pebble accretion with the same scaling of . However, planetesimals can also be scattered by planets and planetary embryos which can reduce the rate of accretion with the exact rate dependent on the assumed disk conditions and the size of the embryos and planetesimals (e.g., S. Shibata et al., 2023), all of which are uncertain.
Alternatively, the final stage of giant formation may be that of giant impact between multiple giants (S. Ginzburg & E. Chiang, 2020). It is not immediately obvious how the giant impact stage would interface with high-eccentricity migration. The proto-hot-Jupiter may have completed its giant impact beyond the iceline and was scattered inward or it may have undergone a few giant impact on its route towards short orbital period. One way to test this is to select warm Jupiters at eccentricities high enough to be considered proto-hot-Jupiters (i.e., constant angular momentum consistent with a circular orbit at a few day orbital period) and verify whether they have elevated water abundance compared to lower eccentricity warm Jupiters. A similar test can be done between warm Jupiters that are spin-orbit aligned vs. misaligned (e.g., J. Zak et al., 2025) although orbital misalignment can also realize from local planet-planet scattering (J. Dong et al. in prep).
@@ 5.3 Suggestion for Ariel
Under a uniform period distribution 232 MCS planets, separated into bins of hot and warm Jupiters, we find that Ariel observations at the Tier 2 level of uncertainties are sufficient to test the effect of high-eccentricity migration of hot Jupiters by way of their elevated water abundance ratio compared to warm Jupiters (Figure 7). Even in the relatively pessimistic scenario (perfect dust trap), and with some additional scatter in planet metallicity, 2- detections are possible with under 252 targets on average, though quite a few more are needed for 3- detections, in part resulting from a limited number of warm Jupiters in the MCS (Figure 9). In other words, the challenge is to sample enough hot and warm Jupiters. On the other hand, in the more optimistic case of no dust trapping, less than 100 targets are needed to achieve even - detections.
@@
Our proposed baseline observational test of comparing hot vs. warm Jupiter water abundance is based on the assumption that warm Jupiters form inside the iceline. Such a test would prefer warm Jupiters that are dynamically cold (i.e., low eccentricity, spin-orbit angle, existence of nearby companions). On the other hand, dynamically hot warm Jupiters (i.e., those with high enough eccentricity to be considered proto-hot-Jupiters) would also be valuable targets: if they also feature elevated water abundance, it would further support the high-eccentricity migration origin of hot Jupiters. We suggest including both dynamically hot and dynamically cold warm Jupiter population in the target sample.
@@ 6 Conclusion
We explored three post-formation pollution mechanisms—planetesimal accretion, pebble accretion, and disk-induced migration—to construct a theoretical prediction of gas giants’ atmospheric water abundance as a function of orbital period. Only pebble accretion in highly metal-heavy disks is able to achieve supersolar metallicity in giant planet atmosphere leaving a detectable signature of pollution with the caveat that a more optimistic treatment of pollution by planetesimal accretion can also achieve supersolar metallicity. If the dominant origin channel of hot Jupiters is high-eccentricity migration while that of warm Jupiters is more dynamically quiescent, we expect a significant upward shift in their atmospheric water abundance compared to warm Jupiters, at a level that is detectable with Ariel. To ensure that it can constrain these formation/migration scenarios, the Ariel Tier 2 transit survey should include hot (10 days) and warm (10 days) Jupiters with at least 20 warm Jupiters beyond 10 days. Probing both dynamically hot and dynamically cold warm Jupiters is recommended for a more complete test of high-eccentricity migration origin of hot Jupiters. We further suggest a comparison of median hot Jupiter water abundance with that of wide-orbit cold Jupiters which can be leveraged to reveal the property of primordial protoplanetary disks such as the existence and the efficacy of the dust trap established by the embedded giant planet(s).